OpenStudy (anonymous):

Find the 9th term of the arithmetic sequence. a7 = -42 and a11 = -28; a9 = ____.

OpenStudy (amistre64):

find the distance from a7 to a11; and divide it by how many steps it takes to get from a7 to a11

OpenStudy (amistre64):

it takes 14 units to get from -42 to -28

OpenStudy (amistre64):

a7 = -42 a8 = -42 + n a9 = -42 + n + n a10 = -42 + n + n + n a11 = -42 + n + n + n +n = -28 -42 +4n = -28 4n = -28 +42 n = 14/4 if I see it right

OpenStudy (amistre64):

a9 = -42 +2(14/4) a9 = -42 +14/2 a9 = -42 + 7 a9 = -35

OpenStudy (anonymous):

Find the 8th term of the arithmetic sequence: 3, 8, 13, 18, ...?

OpenStudy (anonymous):

a11=a7+(11-7)*d <=>-28 = -42 +5d <=> d=7/2 a9 = a7 +(9-7)*d = -42 + 2*7/2 = -35

OpenStudy (amistre64):

a1=3 a2=3+5 = 8 a3=3+2(5) = 13 an=3+(n-1)(5) a8 = 3 + 7(5) = 38

OpenStudy (anonymous):

can you please check my other one please you soo helpful